# Semiclassical Gravity Beyond General Relativity: Insights from Torsion

### v1

@@ 5.3.1 Scale ambiguities
${\Sigma^{M^{2}}}_{c}^{\ ab}$ $=-\frac{1}{8\pi^{2}}\lim_{x^{\prime}\rightarrow x}\hat{\sigma}_{c}^{\ ab}(x,x^{\prime})V(x,x^{\prime})\ln(M^{2}).$ (88)
@@
$+\xi[K^{c}_{\ c(a}\mathring{\nabla}_{b)}v_{0}-K^{c}_{\ (ab)}\mathring{\nabla}_{c}v_{0}$
$+(K^{d}_{\ (ab)}K^{c}_{\ cd}-K^{d}_{\ c(a}K^{c}_{\ b)d})v_{0}$
$+g_{ab}\mathring{\nabla}_{c}(K^{c\ d}_{\ d}v_{0})]\Bigg\}\ln(M^{2}),$ (89)
@@
$+(g^{ab}K^{d}_{\ dc}-\delta^{a}_{\ c}K^{d\ b}_{\ d}+K_{c}^{\ ba}-K^{b\ a}_{\ c})v_{0}\}\ln(M^{2}).$ (90)
Using Eqs. (62) and (63), one can verify that Eq. (5.2) is satisfied for $\Theta^{M^{2}}_{ab}$ and ${\Sigma^{M^{2}}}_{c}^{\ ab}$. In fact, $\mathring{\nabla}^{a}\Theta^{M^{2}}_{ab}$ and ${\Sigma^{M^{2}}}_{c}^{\ ab}$ assume the same form as Eqs. (5.2) and (82), respectively, under the replacement $w\rightarrow w-v_{0}\ln(M^{2})$. Consequently, Eq. (5.2) is automatically satisfied, and the scale ambiguities are consistent with the third axiom.
@@ 5.3.2 Renormalization ambiguities
The renormalization ambiguities, $\Theta_{ab}$ and $\Sigma_{c}^{\ ab}$, can be derived from a renormalization Lagrangian $\mathcal{L}_{\text{Ren}}=\mathcal{L}_{\text{Ren}}[g^{ab},K^{c}_{\ ab}]$ by defining
@@
$\Sigma_{c}^{\ ab}$ $=-\frac{1}{\sqrt{-g}}\frac{\delta(\mathcal{L}_{\text{Ren}}\sqrt{-g})}{\delta K^{c}_{\ ab}}.$ (91)
We emphasize that this Lagrangian must be purely geometrical, as its role is to produce objects that cancel geometrical quantities, and it must remain regular in the flat-spacetime limit. Moreover, by construction, it does not involve nondynamical fields. Consequently, its associated action is invariant under diffeomorphisms [35, 36], ensuring that Eq. (5.2) is automatically satisfied for $\Theta_{ab}$ and $\Sigma_{c}^{\ ab}$. In addition, in four spacetime dimensions, and working in units where $c=\hbar=1$, but without fixing $G$, $\mathcal{L}_{\text{Ren}}$ has units of $\text{length}^{-4}$.
@@
    $*(e^{\alpha_{1}}\wedge\cdots\wedge e^{\alpha_{r}})$ $=\frac{\epsilon^{\alpha_{1}\cdots\alpha_{r}}_{\qquad\ \beta_{r+1}\cdots\beta_{m}}}{(m-r)!}\,e^{\beta_{r+1}}\wedge\cdots\wedge e^{\beta_{m}},$ (94)
@@
$0$ $=$ $\text{d}R^{\mu\nu}+\omega^{\mu}_{\ \rho}\wedge R^{\rho\nu}+\omega^{\nu}_{\ \rho}\wedge R^{\mu\rho},$ (97)
and omitting “topological” terms of the form $\int\text{d}\alpha$, which do not contribute to $\Theta_{ab}$ and $\Sigma_{c}^{\ ab}$, the most general form of $L_{\text{Ren}}$ with dimensions of $\text{length}^{-4}$ is
$L_{\text{Ren}}$ $=\tilde{\alpha}_{1}*(R^{\mu\nu}\wedge e^{\rho}\wedge e^{\sigma})R^{\alpha\beta}\wedge e^{\gamma}\wedge e^{\delta}\epsilon_{\mu\nu\rho\sigma}\epsilon_{\alpha\beta\gamma\delta}$
$+\tilde{\alpha}_{2}*(R^{\mu\nu}\wedge e^{\rho}\wedge e^{\sigma})R_{\mu\nu}\wedge e_{\rho}\wedge e_{\sigma}+\tilde{\alpha}_{3}*(R^{\mu\nu})\wedge R^{\alpha\beta}\epsilon_{\mu\nu\alpha\beta}$
$+\tilde{\alpha}_{4}*(T^{\mu}\wedge T_{\mu})R^{\nu\rho}\wedge e^{\alpha}\wedge e^{\beta}\epsilon_{\nu\rho\alpha\beta}+\tilde{\alpha}_{5}*(T^{\mu}\wedge T_{\mu})R^{\nu\rho}\wedge e_{\nu}\wedge e_{\rho}$
$+\tilde{\alpha}_{6}*(T_{\mu}\wedge T_{\nu})R^{\mu\rho}\wedge e^{\nu}\wedge e_{\rho}+\tilde{\alpha}_{7}*(T^{\mu}\wedge T_{\mu})T^{\nu}\wedge T_{\nu}$
$+\tilde{\alpha}_{8}*(T^{\mu}\wedge T^{\nu})T_{\mu}\wedge T_{\nu}+\tilde{\beta}_{1}R^{\mu\nu}\wedge e_{\mu}\wedge e_{\nu}+\tilde{\beta}_{2}R^{\mu\nu}\wedge e^{\rho}\wedge e^{\sigma}\epsilon_{\mu\nu\rho\sigma}$
$+\tilde{\gamma}e^{\mu}\wedge e^{\nu}\wedge e^{\rho}\wedge e^{\sigma}\epsilon_{\mu\nu\rho\sigma},$ (98)
@@
$+\frac{1}{3}\omega_{\mu\nu}\wedge\omega^{\nu\rho}\wedge\omega_{\rho}^{\ \mu}\Big),$ (100)
@@
$+\frac{1}{3}\omega^{\mu\nu}\wedge\omega^{\rho\kappa}\wedge\omega_{\kappa}^{\ \sigma}\Big)\Big].$ (101)
@@
$\mathcal{L}_{\text{Ren}}$ $=\alpha_{1}R^{2}+\alpha_{2}R_{abcd}R^{abcd}+\alpha_{3}R^{ab}_{\ \ ef}R^{cdef}\epsilon_{abcd}$
$+\alpha_{4}RK_{\epsilon}+\alpha_{5}U^{e}_{\ eabcd}R^{abcd}+\alpha_{6}V_{abcd}R^{abcd}$
$+\alpha_{7}U^{e}_{\ eabcd}K^{fab}K_{f}^{\ cd}+\alpha_{8}U^{(ef)}_{\quad\ abcd}K_{e}^{\ ab}K_{f}^{\ cd}$
@@
$K_{\epsilon}$ $=K^{e}_{\ ab}K_{ecd}\epsilon^{abcd},$ (103)
$U^{ef}_{\ \ abcd}$ $=K^{e}_{\ [ab]}K^{f}_{\ [cd]}-K^{e}_{\ [ac]}K^{f}_{\ [bd]}+K^{e}_{\ [ad]}K^{f}_{\ [bc]},$ (104)
$V^{a}_{\ bcd}$ $=K^{e}_{\ eb}K^{a}_{\ [cd]}+K^{e}_{\ ec}K^{a}_{\ [bd]}-K^{e}_{\ ed}K^{a}_{\ [bc]}$
$-2\left(K^{a}_{\ [be]}K^{e}_{\ [cd]}+K^{a}_{\ [ce]}K^{e}_{\ [bd]}-K^{a}_{\ [de]}K^{e}_{\ [bc]}\right),$ (105)
which satisfy $U^{ef}_{\ \ abcd}=U^{ef}_{\ \ [ab]cd}=U^{ef}_{\ \ ab[cd]}$, $U^{(ef)}_{\quad\ abcd}=U^{(ef)}_{\quad\ cdab}$, and $V^{a}_{\ bcd}=V^{a}_{\ b[cd]}$.
We present the contributions of each term to $\Theta_{ab}$ and $\Sigma_{c}^{\ ab}$ in the following tables, where each row corresponds to a single renormalization term, indicated by its coupling constant.
$\Theta_{ab}$ $\Sigma_{c}^{\ ab}$
@@
$\alpha_{2}$ $\begin{aligned} &-2\Big\{2R_{acde}R_{b}^{\ cde}-\frac{1}{2}R_{cdef}R^{cdef}g_{ab}\\ &+4\mathring{\nabla}_{d}\mathring{\nabla}_{c}R_{(a\ \ b)}^{\ \ \ c\ \ d}\\ &+4\mathring{\nabla}_{e}(R_{(a\ \ b)}^{\ \ \ c\ \ d}K^{e}_{\ cd}+R_{(a}^{\ \ cde}K_{b)cd})\Big\}\end{aligned}$ $\begin{aligned} &-4\{-\mathring{\nabla}_{d}R^{adb}_{\quad c}+R^{adb}_{\quad e}K^{e}_{\ dc}+R^{ad\ e}_{\ \ \ c}K^{b}_{\ de}\}\end{aligned}$
$\alpha_{3}$ $\begin{aligned} &-2\{4R_{(a|}^{\ \ \ cf\ell}R^{de}_{\ \ f\ell}\epsilon_{|b)cde}\\ &-R_{cd\ell m}R_{ef}^{\ \ \ \ell m}\epsilon^{cdef}g_{ab}\\ &+4\mathring{\nabla}_{e}(R^{f\ell\ \ d}_{\ \ (a|}K^{e}_{\ cd}+R^{f\ell de}K_{(a|cd})\epsilon_{|b)\ f\ell}^{\quad c}\\ &+4\mathring{\nabla}_{d}\mathring{\nabla}_{c}R^{ef\ \ d}_{\ \ (a}\epsilon_{b)\ ef}^{\ \ \ c}\}\end{aligned}$ $\begin{aligned} &-4\{-\mathring{\nabla}_{d}R^{efb}_{\quad c}\epsilon^{ad}_{\ \ ef}\\ &+(R^{f\ell b}_{\quad e}K^{e}_{\ dc}+R^{f\ell\ e}_{\ \ c}K^{b}_{\ de})\epsilon^{ad}_{\ \ f\ell}\}\end{aligned}$
@@
$\alpha_{5}$ $\begin{aligned} &-2\Big\{2U^{f}_{\ f(a|cde}R_{|b)}^{\quad cde}+U^{f}_{\ fcde(a|}R^{cde}_{\quad|b)}\\ &-U_{(ab)cdef}R^{cdef}g_{ab}-\frac{1}{2}U^{\ell}_{\ \ell cdef}R^{cdef}g_{ab}\\ &+2[\mathring{\nabla}_{d}\mathring{\nabla}_{c}U^{e\ \ \ c\ \ d}_{\ e(a\ b)}\\ &+\mathring{\nabla}_{e}(U^{f\ \ \ c\ \ d}_{\ f(a\ b)}+U^{f\ \ \ cde}_{\ f(a}K_{b)cd})]\Big\}\end{aligned}$ $\begin{aligned} &-2\{-\mathring{\nabla}_{d}U^{e\ adb}_{\ e\quad\ c}+U^{f\ adb}_{\ f\quad\ e}K^{e}_{\ dc}\\ &+U^{f\ ad\ e}_{\ f\ \ c}K^{b}_{\ de}\\ &+2g_{cf}(R^{a[b|de}+R^{dea[b|}+R^{ae[b|d}\\ &\quad-R^{[b|ead}-R^{ad[b|e}+R^{[b|dae})K^{|f]}_{\quad de}\}\end{aligned}$
@@
$\alpha_{6}$ $\begin{aligned} &-2\Big\{-V^{c\ \ d}_{\ (a|\ e}R_{|b)cd}^{\quad\ \ e}+V^{cd}_{\ \ (a|e}R_{cd|b)}^{\quad\ \ e}\\ &-\frac{1}{2}V_{cdef}R^{cdef}g_{ab}\\ &+\mathring{\nabla}_{d}\mathring{\nabla}_{c}(V_{(a\ b)}^{\ \ c\ \ d}+V^{c\ \ d}_{\ (a\ \ b)})\\ &+\mathring{\nabla}_{e}[(V_{(a\ b)}^{\ \ c\ \ d}+V^{c\ \ d}_{\ (a\ \ b)})K^{e}_{\ cd}\\ &+(V_{(a}^{\ cde}-V^{c\ de}_{\ (a})K_{b)cd}]\Big\}\end{aligned}$ $\begin{aligned} &-\Big\{2[-\mathring{\nabla}_{d}V^{[ad]b}_{c}+V^{[ad]b}_{e}K^{e}_{\ dc}\\ &+V^{[ad]\ e}_{\quad\ \ c}K^{b}_{\ de}]+\Big[g^{ab}\Big(\frac{1}{2}R_{cd}^{\ \ \ ef}-R_{d\ \ c}^{\ \ [e\ f]}\Big)\\ &-\delta^{a}_{\ c}\Big(\frac{1}{2}R^{b\ ef}_{\ d}-R_{d}^{\ [e|b|f]}\Big)\Big]K^{d}_{\ ef}+\Big[R^{ab}_{\ \ cd}\\ &-R^{[a\ b]}_{\ \ c\ \ d}-\frac{1}{2}\Big(R^{[a\ b]}_{\ \ d\ \ c}+R_{cd}^{\ \ \ ab}\Big)\Big]K^{d\ e}_{\ e}\\ &+(R^{[a|\ de}_{\ \ \ c}-2R_{c}^{\ d[a|e})K^{|b]}_{\quad[de]}\\ &-(R^{abde}+2R^{[a|d|b]e})K_{c[de]}\\ &-(R^{de[a}_{\quad\ [c}+2R^{[a|d\ \ e}_{\quad[c})K^{|b]}_{\quad e]d}\\ &-(R^{deab}+2R^{[a|d|b]e})K_{d[ec]}\Big\}\end{aligned}$
$\alpha_{7}$ $\begin{aligned} &-2\Big\{-U_{(ab)cdef}K^{\ell cd}K_{\ell}^{\ ef}\\ &-U^{\ell}_{\ \ell cdef}K_{(a}^{\ \ cd}K_{b)}^{\ \ ef}\\ &+U^{f}_{\ f(a|cde}(K^{\ell\ \ c}_{\ |b)}-K^{\ell c}_{\ \ |b)})K_{\ell}^{\ de}\\ &+U^{f}_{\ fcd(a|e}K^{\ell cd}(K_{\ell|b)}^{\quad e}-K_{\ell\ |b)}^{\ e})\\ &-\frac{1}{2}U^{\ell}_{\ \ell cdef}K^{mcd}K_{m}^{\ \ ef}g_{ab}\Big\}\end{aligned}$ $6g_{c\ell}U^{f\ \ a[b|de}_{\ f}K^{|\ell]}_{\quad de}$

### v2

@@ 5.3.1 Scale ambiguities
${\Sigma^{M^{2}}}_{c}^{\phantom{c}ab}$ $=-\frac{1}{8\pi^{2}}\lim_{x^{\prime}\rightarrow x}\hat{\sigma}_{c}^{\phantom{c}ab}(x,x^{\prime})V(x,x^{\prime})\ln(M^{2}).$ (88)
@@
$+\xi[K^{c}_{\phantom{c}c(a}\mathring{\nabla}_{b)}v_{0}-K^{c}_{\phantom{c}(ab)}\mathring{\nabla}_{c}v_{0}$
$+(K^{d}_{\phantom{d}(ab)}K^{c}_{\phantom{c}cd}-K^{d}_{\phantom{d}c(a}K^{c}_{\phantom{c}b)d})v_{0}$
$+g_{ab}\mathring{\nabla}_{c}(K^{c\phantom{d}d}_{\phantom{c}d}v_{0})]\Bigg\}\ln(M^{2}),$ (89)
@@
$+(g^{ab}K^{d}_{\phantom{d}dc}-\delta^{a}_{\phantom{a}c}K^{d\phantom{d}b}_{\phantom{d}d}+K_{c}^{\phantom{c}ba}-K^{b\phantom{c}a}_{\phantom{b}c})v_{0}\}\ln(M^{2}).$ (90)
Using Eqs. (62) and (63), one can verify that Eq. (5.2) is satisfied for $\Theta^{M^{2}}_{ab}$ and ${\Sigma^{M^{2}}}_{c}^{\phantom{c}ab}$. In fact, $\mathring{\nabla}^{a}\Theta^{M^{2}}_{ab}$ and ${\Sigma^{M^{2}}}_{c}^{\phantom{c}ab}$ assume the same form as Eqs. (5.2) and (82), respectively, under the replacement $w\rightarrow w-v_{0}\ln(M^{2})$. Consequently, Eq. (5.2) is automatically satisfied, and the scale ambiguities are consistent with the third axiom.
@@ 5.3.2 Renormalization ambiguities
The renormalization ambiguities, $\Theta_{ab}$ and $\Sigma_{c}^{\phantom{c}ab}$, can be derived from a renormalization Lagrangian $\mathcal{L}_{\text{Ren}}=\mathcal{L}_{\text{Ren}}[g^{ab},K^{c}_{\phantom{c}ab}]$ by defining
@@
$\Sigma_{c}^{\phantom{c}ab}$ $=-\frac{1}{\sqrt{-g}}\frac{\delta(\mathcal{L}_{\text{Ren}}\sqrt{-g})}{\delta K^{c}_{\phantom{c}ab}}.$ (91)
We emphasize that this Lagrangian must be purely geometrical, as its role is to produce objects that cancel geometrical quantities, and it must remain regular in the flat-spacetime limit. Moreover, by construction, it does not involve nondynamical fields. Consequently, its associated action is invariant under diffeomorphisms [35, 36], ensuring that Eq. (5.2) is automatically satisfied for $\Theta_{ab}$ and $\Sigma_{c}^{\phantom{c}ab}$. In addition, in four spacetime dimensions, and working in units where $c=\hbar=1$, but without fixing $G$, $\mathcal{L}_{\text{Ren}}$ has units of $\text{length}^{-4}$.
@@
[…]
@@
$0$ $=$ $\text{d}R^{\mu\nu}+\omega^{\mu}_{\phantom{\mu}\rho}\wedge R^{\rho\nu}+\omega^{\nu}_{\phantom{\nu}\rho}\wedge R^{\mu\rho},$ (97)
and omitting “topological” terms of the form $\int\text{d}\alpha$, which do not contribute to $\Theta_{ab}$ and $\Sigma_{c}^{\phantom{c}ab}$, the most general form of $L_{\text{Ren}}$ with dimensions of $\text{length}^{-4}$ is
$L_{\text{Ren}}$ $=\tilde{\alpha}_{1}*(R^{\mu\nu}\wedge e^{\rho}\wedge e^{\sigma})R^{\alpha\beta}\wedge e^{\gamma}\wedge e^{\delta}\epsilon_{\mu\nu\rho\sigma}\epsilon_{\alpha\beta\gamma\delta}+\tilde{\alpha}_{2}*(R^{\mu\nu}\wedge e^{\rho}\wedge e^{\sigma})R_{\mu\nu}\wedge e_{\rho}\wedge e_{\sigma}$
$+\tilde{\alpha}_{3}*(R^{\mu\nu})\wedge R^{\alpha\beta}\epsilon_{\mu\nu\alpha\beta}+\tilde{\alpha}_{4}*(T^{\mu}\wedge T_{\mu})R^{\nu\rho}\wedge e^{\alpha}\wedge e^{\beta}\epsilon_{\nu\rho\alpha\beta}+\tilde{\alpha}_{5}*(T^{\mu}\wedge T_{\mu})R^{\nu\rho}\wedge e_{\nu}\wedge e_{\rho}$
$+\tilde{\alpha}_{6}*(T_{\mu}\wedge T_{\nu})R^{\mu\rho}\wedge e^{\nu}\wedge e_{\rho}+\tilde{\alpha}_{7}*(T^{\mu}\wedge T_{\mu})T^{\nu}\wedge T_{\nu}+\tilde{\alpha}_{8}*(T^{\mu}\wedge T^{\nu})T_{\mu}\wedge T_{\nu}+\tilde{\beta}_{1}R^{\mu\nu}\wedge e_{\mu}\wedge e_{\nu}$
$+\tilde{\beta}_{2}R^{\mu\nu}\wedge e^{\rho}\wedge e^{\sigma}\epsilon_{\mu\nu\rho\sigma}+\tilde{\gamma}e^{\mu}\wedge e^{\nu}\wedge e^{\rho}\wedge e^{\sigma}\epsilon_{\mu\nu\rho\sigma},$ (98)
@@
$+\frac{1}{3}\omega_{\mu\nu}\wedge\omega^{\nu\rho}\wedge\omega_{\rho}^{\phantom{\rho}\mu}\Big),$ (100)
@@
$+\frac{1}{3}\omega^{\mu\nu}\wedge\omega^{\rho\kappa}\wedge\omega_{\kappa}^{\phantom{\kappa}\sigma}\Big)\Big].$ (101)
@@
$\mathcal{L}_{\text{Ren}}$ $=\alpha_{1}R^{2}+\alpha_{2}R_{abcd}R^{abcd}+\alpha_{3}R^{ab}_{\phantom{ab}ef}R^{cdef}\epsilon_{abcd}$
$+\alpha_{4}RK_{\epsilon}+\alpha_{5}U^{e}_{\phantom{e}eabcd}R^{abcd}+\alpha_{6}V_{abcd}R^{abcd}$
$+\alpha_{7}U^{e}_{\phantom{e}eabcd}K^{fab}K_{f}^{\phantom{f}cd}+\alpha_{8}U^{(ef)}_{\phantom{(ef)}\ abcd}K_{e}^{\phantom{e}ab}K_{f}^{\phantom{f}cd}$
@@
$K_{\epsilon}$ $=K^{e}_{\phantom{e}ab}K_{ecd}\epsilon^{abcd},$ (103)
$U^{ef}_{\phantom{ef}abcd}$ $=K^{e}_{\phantom{e}[ab]}K^{f}_{\phantom{f}[cd]}-K^{e}_{\phantom{e}[ac]}K^{f}_{\phantom{f}[bd]}+K^{e}_{\phantom{e}[ad]}K^{f}_{\phantom{f}[bc]},$ (104)
$V^{a}_{\phantom{a}bcd}$ $=K^{e}_{\phantom{e}eb}K^{a}_{\phantom{a}[cd]}+K^{e}_{\phantom{e}ec}K^{a}_{\phantom{a}[bd]}-K^{e}_{\phantom{e}ed}K^{a}_{\phantom{a}[bc]}$
$-2\left(K^{a}_{\phantom{a}[be]}K^{e}_{\phantom{e}[cd]}+K^{a}_{\phantom{a}[ce]}K^{e}_{\phantom{e}[bd]}-K^{a}_{\phantom{a}[de]}K^{e}_{\phantom{e}[bc]}\right),$ (105)
which satisfy $U^{ef}_{\phantom{ef}abcd}=U^{ef}_{\phantom{ef}[ab]cd}=U^{ef}_{\phantom{ef}ab[cd]}$, $U^{(ef)}_{\phantom{(ef)}\ abcd}=U^{(ef)}_{\phantom{(ef)}\ cdab}$, and $V^{a}_{\phantom{a}bcd}=V^{a}_{\phantom{a}b[cd]}$.
We present the contributions of each term to $\Theta_{ab}$ and $\Sigma_{c}^{\phantom{c}ab}$ in the following tables, where each row corresponds to a single renormalization term, indicated by its coupling constant.
$\Theta_{ab}$ $\Sigma_{c}^{\phantom{c}ab}$
@@
$\alpha_{2}$ $\begin{aligned} &-2\Big\{2R_{acde}R_{b}^{\phantom{b}cde}-\frac{1}{2}R_{cdef}R^{cdef}g_{ab}+4\mathring{\nabla}_{d}\mathring{\nabla}_{c}R_{(a\phantom{c}b)}^{\phantom{(a}c\phantom{b)}d}\\ &\qquad+4\mathring{\nabla}_{e}(R_{(a\phantom{c}b)}^{\phantom{(a}c\phantom{b)}d}K^{e}_{\phantom{e}cd}+R_{(a}^{\phantom{(a}cde}K_{b)cd})\Big\}\end{aligned}$ $\begin{aligned} &-4\{-\mathring{\nabla}_{d}R^{adb}_{\phantom{adb}c}+R^{adb}_{\phantom{adb}e}K^{e}_{\phantom{e}dc}+R^{ad\phantom{c}e}_{\phantom{ad}c}K^{b}_{\phantom{b}de}\}\end{aligned}$
$\alpha_{3}$ $\begin{aligned} &-2\{4R_{(a|}^{\phantom{(a|}cf\ell}R^{de}_{\phantom{de}f\ell}\epsilon_{|b)cde}-R_{cd\ell m}R_{ef}^{\phantom{ef}\ell m}\epsilon^{cdef}g_{ab}\\ &\qquad+4\mathring{\nabla}_{d}\mathring{\nabla}_{c}R^{ef\phantom{(a}d}_{\phantom{ef}(a}\epsilon_{b)\phantom{c}ef}^{\phantom{b)}c}\\ &\qquad+4\mathring{\nabla}_{e}(R^{f\ell\phantom{(a|}d}_{\phantom{ef}(a|}K^{e}_{\phantom{e}cd}+R^{f\ell de}K_{(a|cd})\epsilon_{|b)\phantom{c}f\ell}^{\phantom{|b)}c}\}\end{aligned}$ $\begin{aligned} &-4\{-\mathring{\nabla}_{d}R^{efb}_{\phantom{efb}c}\epsilon^{ad}_{\phantom{ad}ef}+(R^{f\ell b}_{\phantom{f\ell b}e}K^{e}_{\phantom{e}dc}+R^{f\ell\phantom{c}e}_{\phantom{f\ell}c}K^{b}_{\phantom{b}de})\epsilon^{ad}_{\phantom{ad}f\ell}\}\end{aligned}$
@@
$\alpha_{5}$ $\begin{aligned} &-2\Big\{2U^{f}_{\phantom{f}f(a|cde}R_{|b)}^{\phantom{|b)}cde}+U^{f}_{\phantom{f}fcde(a|}R^{cde}_{\phantom{cde}|b)}\\ &\qquad-U_{(ab)cdef}R^{cdef}g_{ab}-\frac{1}{2}U^{\ell}_{\phantom{\ell}\ell cdef}R^{cdef}g_{ab}\\ &\qquad+2[\mathring{\nabla}_{d}\mathring{\nabla}_{c}U^{e\phantom{e(a}c\phantom{b)}d}_{\phantom{e}e(a\phantom{c}b)}\\ &\qquad\quad\ \ +\mathring{\nabla}_{e}(U^{f\phantom{f(a}c\phantom{b)}d}_{\phantom{f}f(a\phantom{c}b)}K^{e}_{\phantom{e}cd}+U^{f\phantom{f(a}cde}_{\phantom{f}f(a}K_{b)cd})]\Big\}\end{aligned}$ $\begin{aligned} &-2\{-\mathring{\nabla}_{d}U^{e\phantom{e}adb}_{\phantom{e}e\phantom{adb}c}+U^{f\phantom{f}adb}_{\phantom{f}f\phantom{adb}e}K^{e}_{\phantom{e}dc}+U^{f\phantom{f}ad\phantom{c}e}_{\phantom{f}f\phantom{ad}c}K^{b}_{\phantom{b}de}\\ &\qquad+2g_{cf}(R^{a[b|de}+R^{dea[b|}+R^{ae[b|d}\\ &\qquad\qquad\quad-R^{[b|ead}-R^{ad[b|e}+R^{[b|dae})K^{|f]}_{\phantom{|f]}de}\}\end{aligned}$
@@
$\alpha_{6}$ $\begin{aligned} &-2\Big\{-V^{c\phantom{(a|}d}_{\phantom{c}(a|\phantom{d}e}R_{|b)cd}^{\phantom{|b)cd}e}+V^{cd}_{\phantom{cd}(a|e}R_{cd|b)}^{\phantom{cd|b)}e}\\ &\qquad-\frac{1}{2}V_{cdef}R^{cdef}g_{ab}+\mathring{\nabla}_{d}\mathring{\nabla}_{c}(V_{(a\phantom{c}b)}^{\phantom{(a}c\phantom{b)}d}+V^{c\phantom{(a}d}_{\phantom{c}(a\phantom{d}b)})\\ &\qquad+\mathring{\nabla}_{e}[(V_{(a\phantom{c}b)}^{\phantom{(a}c\phantom{b)}d}+V^{c\phantom{(a}d}_{\phantom{c}(a\phantom{d}b)})K^{e}_{\phantom{e}cd}\\ &\qquad+(V_{(a}^{\phantom{(a}cde}-V^{c\phantom{(a}de}_{\phantom{c}(a})K_{b)cd}]\Big\}\end{aligned}$ $\begin{aligned} &-\Big\{2[-\mathring{\nabla}_{d}V^{[ad]b}_{\phantom{[ad]b}c}+V^{[ad]b}_{\phantom{[ad]b}e}K^{e}_{\phantom{e}dc}+V^{[ad]\phantom{c}e}_{\phantom{[ad]}c}K^{b}_{\phantom{b}de}]\\ &\quad+\Big[g^{ab}\Big(\frac{1}{2}R_{cd}^{\phantom{cd}ef}-R_{d\phantom{[e}c}^{\phantom{d}[e\phantom{c}f]}\Big)\\ &\qquad-\delta^{a}_{\phantom{a}c}\Big(\frac{1}{2}R^{b\phantom{d}ef}_{\phantom{b}d}-R_{d}^{\phantom{d}[e|b|f]}\Big)\Big]K^{d}_{\phantom{d}ef}\\ &\quad+\Big[-\frac{1}{2}\Big(R^{[a\phantom{d}b]}_{\phantom{[a}d\phantom{b]}c}+R_{cd}^{\phantom{cd}ab}\Big)+R^{ab}_{\phantom{ab}cd}-R^{[a\phantom{c}b]}_{\phantom{[a}c\phantom{b]}d}\Big]K^{d\phantom{e}e}_{\phantom{d}e}\\ &\quad+(R^{[a|\phantom{c}de}_{\phantom{[a|}c}-2R_{c}^{\phantom{c}d[a|e})K^{|b]}_{\phantom{|b]}[de]}\\ &\qquad-(R^{abde}+2R^{[a|d|b]e})K_{c[de]}\\ &\quad-(R^{de[a}_{\phantom{de[a}[c}+2R^{[a|d\phantom{[c}e}_{\phantom{[a|d}[c})K^{|b]}_{\phantom{|b]}e]d}\\ &\qquad-(R^{deab}+2R^{[a|d|b]e})K_{d[ec]}\Big\}\end{aligned}$
$\alpha_{7}$ $\begin{aligned} &-2\Big\{-U_{(ab)cdef}K^{\ell cd}K_{\ell}^{\phantom{\ell}ef}-U^{\ell}_{\phantom{\ell}\ell cdef}K_{(a}^{\phantom{(a}cd}K_{b)}^{\phantom{b)}ef}\\ &\qquad+U^{f}_{\phantom{f}f(a|cde}(K^{\ell\phantom{|b)}c}_{\phantom{\ell}|b)}-K^{\ell c}_{\phantom{\ell c}|b)})K_{\ell}^{\phantom{\ell}de}\\ &\qquad+U^{f}_{\phantom{f}fcd(a|e}K^{\ell cd}(K_{\ell|b)}^{\phantom{\ell|b)}e}-K_{\ell\phantom{e}|b)}^{\phantom{\ell}e})\\ &\qquad-\frac{1}{2}U^{\ell}_{\phantom{\ell}\ell cdef}K^{mcd}K_{m}^{\phantom{m}ef}g_{ab}\Big\}\end{aligned}$ $6g_{c\ell}U^{f\phantom{f}a[b|de}_{\phantom{f}f}K^{|\ell]}_{\phantom{|\ell]}de}$